### v1

# Coherence and Entanglement in a Quantum Spacetime

Abstract
We investigate the emergence of quantum coherence and quantum correlations in a two-particle system with deformed symmetries arising from the quantum nature of spacetime. We demonstrate that the deformation of energy-momentum composition induces a momentum-dependent interaction that counteracts the decoherence effects described by the Lindblad equation in quantum spacetime. This interplay leads to the formation of coherence, entanglement and other correlations, which we quantify using concurrence, the $l_{1}$-norm of coherence, quantum discord and Local Quantum Fisher Information. Our analysis reveals that while the openness of quantum spacetime ultimately degrades entanglement, it also facilitates the creation and preservation of both classical and quantum correlations.
I Introduction
Over the past few decades, significant theoretical efforts have been devoted to formulating a quantum theory of gravity. Various proposals have emerged, including string theory Polchinski:1998rq , loop quantum gravity Ashtekar:2021kfp , causal dynamical triangulation Loll:2019rdj , and asymptotically safe gravity Eichhorn:2018yfc , among others. One of the most intriguing aspects of these approaches is the possibility that Lorentz’s symmetry may not be a fundamental symmetry of nature. At the Planck scale, this could imply either the absence of local equivalent frames or that Lorentz symmetry is an intermediate step toward a more fundamental symmetry. The latter scenario suggests that the principle of relativity could be compatible with the existence of a fundamental length scale Amelino-Camelia:2000stu , a possibility currently under experimental investigation Amelino-Camelia:2008aez ; Addazi:2021xuf ; AlvesBatista:2023wqm .
A quantum spacetime with nontrivial features can leave possibly observable imprints on quantum mechanical systems. These include violations of the equivalence principle Wagner:2023fmb and the emergence of fractal dimensions Varao:2024eig . Moreover, quantum spacetime may act as an open environment, therefore leading to fundamental decoherence in quantum systems Breuer:2008rh ; Streltsov:2016iow . This phenomenon can be associated with a violation of the relativity principle, where Lorentz (or Galilean) transformations cease to be symmetries Petruzziello:2020wkd . Moreover, even when the relativity principle is preserved at the Planck scale, decoherence effects can still arise as a consequence of a deformed time evolution operator. This produces a modification of the evolution equation of quantum systems described by a Lindbladian operator Arzano:2022nlo . While this has been established for single-particle states, the effects on multiparticle systems have not been addressed. In this work, we investigate this, focusing on two-particle systems.
Modified spacetime symmetry transformations Lukierski:1991pn ; Lukierski:1993wxa can alter conservation laws Majid:1994cy ; Amelino-Camelia:2000stu ; Borowiec:2009vb , thus giving rise to the concept of relative locality Amelino-Camelia:2011lvm ; Gubitosi:2011hgc . One intriguing possibility is the modification of energy conservation, which, at the quantum level, implies that the Hamiltonian of a two-particle system could acquire additional nontrivial contributions leading to decoherence, which do not have relevant implications in the analysis of single-particle decoherence. In addition to decoherence effects, depending on the nature of the additional terms in the Hamiltonian, correlations between particles can emerge and potentially compete with decoherence effects. For instance, recent studies have explored the formation of entanglement due to gravitational interactions between massive states in spatial superpositions Bose:2017nin ; Marletto:2017kzi , offering a potential window into quantum gravity. Entanglement, a quintessentially quantum correlation with no classical counterpart Horodecki:2009zz , is one such possibility. Beyond entanglement, quantum discord represents the first historical measure of quantum correlations PhysRevLett.88.017901 , defined as the difference between mutual information and classical correlations.
In this paper, we examine quantum correlation through the evolution of three key quantifiers: the concurrence as a measure of entanglement Wootters:1997id , the $l_{1}$-norm of coherence to track decoherence effects Baumgratz:2013ecx , the quantum discord as defined in PhysRevLett.88.017901 and the local quantum Fisher information slaoui2019comparative ; Silva:2025vmu ; kim2018characterizing ; Dahbi:2022uyi ; Lian:2021xjw ; Yurischev:2023jrd .
The paper is organized as follows. In Section II, we revisit the formalization of deformed symmetries within a Hopf algebra model and the derivation of a Lindblad equation. In Section III, we construct a model of two quantum particles with two energy levels, incorporating an interaction term arising from the deformed composition of energy. In Section IV, we analyze quantum correlations in this model. In Section V, we explore how the deformed Hamiltonian and the Lindblad term influence the interplay between these correlations. Finally, in Section VI, we present our conclusions. Throughout the paper, we use natural units where $\hbar=c=1$, implying that the Planck energy is the inverse of the Planck length.
II Deformation of Spacetime Symmetries
A generic feature of non-commutative spacetimes is that the description of free particles is generalized to non-commutative ordered exponentials Majid:1994cy . As a result, quantum states in the momentum representation exhibit novel properties, such as the deformation of energy and momentum conservation laws. Specifically, the operator that calculates the total momentum of a two-particle state $|g\rangle\otimes|h\rangle$ is given by a coproduct $\Delta P_{\mu}$, which, in general, does not follow the usual linear rule for composing the energy and momentum of the two-particle system. These modifications are suppressed by a noncommutativity parameter, expected to be of the order of the quantum gravity scale. An interesting aspect of this approach is that it admits a nonrelativistic limit that retains traces of the quantum gravity scale. Working with the classical basis of $\kappa$-Poincaré algebra, i.e, the one in which the dispersion relation is unmodified, in the nonrelativistic (Galilean) limit the modified conservation law takes the form:
$\Delta P_{0}$ $=P_{0}\otimes\mathbb{1}+\mathbb{1}\otimes P_{0}-i\ell\sum_{i}P_{i}\otimes P_{i},$ (1)
$\Delta P_{i}$ $=P_{i}\otimes\mathbb{1}+\mathbb{1}\otimes P_{i},$ (2)
where $\ell^{-1}$ is the quantum gravity energy scale, often denoted as $\kappa$. This operation can be used to define the inverse of the translation group element via the antipode map $S(P)$, where ($\mathbf{P}^{2}=\sum_{i}(P_{i})^{2}$):
$S(P_{0})$ $=-P_{0}-i\ell\mathbf{P}^{2},$ (3)
$S(P_{i})$ $=-P_{i}+\ell P_{i}P_{0}.$ (4)
The form of the coproduct operators leads to a modified, non-Abelian conservation law for energy and momentum. When applied to a composite system $|g\rangle\otimes|h\rangle$, the eigenvalues are given by:
$g_{0}\oplus h_{0}$ $=g_{0}+h_{0}+\ell\vec{g}\cdot\vec{h},$ (5)
$g_{i}\oplus h_{i}$ $=g_{i}+h_{i},$ (6)
where $(g_{0},g_{i})$ represents the energy and momentum (in Cartesian coordinates) of the single-particle state $|g\rangle$ (and similarly for $h$).
II.1 Lindblad Equation
The presence of deformed symmetries induces a deformation of the commutator, given by the adjoint action Arzano:2022nlo . This action selects a representative of the vector space ${\cal R}_{G}$, spanned by the generators of the symmetry algebra $G\in{\cal R}_{G}$, and acts on the quantum system ${\cal A}_{\cal H}$ defined on a Hilbert space ${\cal H}$ as:
$\text{ad}_{\clubsuit}\spadesuit\,:\,{\cal R}_{G}\times{\cal A}_{\cal H}\rightarrow{\cal A}_{\cal H},\quad\text{ad}_{G}O=(\text{id}\otimes S)\Delta G\diamond O,$ (7)
where id is the identity operator, $\Delta G$ is the coproduct operator, $S$ is the antipode map, and $\diamond$ is the operator defined as:
$(a\otimes b)\diamond O\doteq aOb.$ (8)
When $\Delta G=\sum_{i}a_{i}\otimes b_{i}$, then $(\text{id}\otimes S)\Delta G=\sum_{i}a_{i}\otimes S(b_{i})$, which implies that $\text{ad}_{G}O=\sum_{i}a_{i}OS(b_{i})$. If $O=\rho$ is the density matrix of a quantum system and $\Delta G=P_{0}\otimes\mathbb{1}+\mathbb{1}\otimes P_{0}$, then $\text{ad}_{P_{0}}\rho=[P_{0},\rho]$. This means that the evolution equation of a system with Galilean invariance can be written as $i\partial_{t}\rho=\text{ad}_{P_{0}}\rho$. For a quantum spacetime with deformed symmetries, the generalization of this evolution equation is:
$i\partial_{t}\rho=\frac{1}{2}\left\{\text{ad}_{P_{0}}\rho-\left[\text{ad}_{P_{0}}\rho\right]^{\dagger}\right\}.$ (9)
A straightforward calculation based on the above definition leads to a Lindblad equation for a quantum system with deformed symmetries:
$\partial_{t}\rho=-i[P_{0},\rho]-\frac{\ell}{2}\left(\mathbf{P}^{2}\rho+\rho\mathbf{P}^{2}-2\sum_{i}P_{i}\rho P_{i}\right).$ (10)
The presence of a Lindblad equation implies that a quantum spacetime acts as a source of fundamental decoherence in quantum systems. The case of single particles has been studied in Arzano:2022nlo , where the decoherence time was estimated. The two-particle case introduces new results, which will be analyzed in the next section.
III Two-Particle Model
The modification of the energy conservation law, given by Eq. (5), leads to the definition of the Hamiltonian for a system of two free particles as:
$H=H_{0}\otimes\mathbb{1}+\mathbb{1}\otimes H_{0}+\ell\sum_{i}P_{i}\otimes P_{i},$ (11)
where $H_{0}$ is the Hamiltonian of a free particle. Notably, the new term in (11) resembles a momentum-dependent potential that may induce correlations between the particles.
In the following, we will analyze the case of two interacting particles, each occupying a defined energy state.
III.1 Two-Energy States
To estimate these effects, we consider a simple case of two particles, each modeled as a two-level system. Individually, each particle can occupy energy states with energies $E_{(0)}=\dfrac{\epsilon-\omega}{2}$ or $E_{(1)}=\dfrac{\epsilon+\omega}{2}$, where $\epsilon>\omega$ ensures positive energies. If the energy composition were undeformed, the Hamiltonian of this system would be:
$H_{\text{und}}=\epsilon\mathbb{1}\otimes\mathbb{1}+\frac{\omega}{2}(\sigma_{z}\otimes\mathbb{1}+\mathbb{1}\otimes\sigma_{z}),$ (12)
where $\sigma_{z}$ is the $z$-Pauli matrix. The two-particle state would then be described by two indices $|ij\rangle$, where $(ij)=\{0,1\}$ labels whether the first or second particle is in the ground or excited state.
Now, suppose the particles are in a superposition of momenta. Typically, this would not affect the system’s description, but since the inner product of momenta influences the energy of the composite system, the direction of each momentum becomes important. For simplicity, we consider a one-dimensional case, which suffices for our purposes. Here, each particle can move in either the positive or negative direction, meaning the composite system is described by a state $|ij\alpha\beta\rangle$, where $(\alpha,\beta)=\{+,-\}$ labels the direction of the first or second particle.
Since the particles are well-localized in momentum representation, the eigenvectors of direction for each particle can be written as a linear combination of energy eigenstates:
|±$|\pm\rangle=\frac{1}{\sqrt{2}}(|0\rangle\pm|1\rangle).$ (13)
This approximation is reasonable, as it can be verified by combining energy eigenfunctions of a 1D harmonic oscillator in the momentum representation, showing that this combination is concentrated in the positive or negative branch of momentum.
From the undeformed dispersion relation (since we are at the classical basis), we construct an operator that gives the modulus of the momentum of a two-level state, sharing the same energy eigenstates but with different energies for each level:
$|P|=\text{diag}\left(\sqrt{m(\epsilon+\omega)},\sqrt{m(\epsilon-\omega)}\right).$ (14)
To account for the momentum direction, we note that the $x$-Pauli matrix has eigenstates given by Eq. (13). Thus, the Hamiltonian describing the evolution of this system is:
${\cal H}$ $=\epsilon(\mathbb{1})^{4}+\dfrac{\omega}{2}\left(\sigma_{z}\otimes\mathbb{1}\otimes\mathbb{1}\otimes\mathbb{1}+\mathbb{1}\otimes\sigma_{z}\otimes\mathbb{1}\otimes\mathbb{1}\right)$
$\quad+\ell|P|\otimes|P|\otimes\sigma_{x}\otimes\sigma_{x},$ (15)
where $(\mathbb{1})^{4}=\mathbb{1}\otimes\mathbb{1}\otimes\mathbb{1}\otimes\mathbb{1}$, and $\mathbb{1}$ is the $2\times 2$ identity matrix. This Hamiltonian is a $16\times 16$ matrix, which can be reduced to four $4\times 4$ block matrices as follows:
${\cal H}=\begin{pmatrix}H_{11}&0&0&0\\ 0&H_{-11}&0&0\\ 0&0&H_{1-1}&0\\ 0&0&0&H_{-1-1}\end{pmatrix},$ (16)
where
$H_{ab}=\begin{pmatrix}\epsilon+\frac{a+b}{2}\omega&0&0&\ell m\sqrt{(\epsilon+a\omega)(\epsilon+b\omega)}\\ 0&\epsilon+\frac{a+b}{2}\omega&\ell m\sqrt{(\epsilon+a\omega)(\epsilon+b\omega)}&0\\ 0&\ell m\sqrt{(\epsilon+a\omega)(\epsilon+b\omega)}&\epsilon+\frac{a+b}{2}\omega&0\\ \ell m\sqrt{(\epsilon+a\omega)(\epsilon+b\omega)}&0&0&\epsilon+\frac{a+b}{2}\omega\end{pmatrix},$ (17)
with $(a,b)=\{-1,1\}$ determining the directions of the particles. Since the evolution of this system involves Lindblad operators, we need to describe the momentum of the composite system. Given that the Lindblad operators are multiplied by $\ell$ in (10) and are expected to contribute only slightly, the momentum composition remains undeformed. The operator representing the sum of the momenta of the two particles is:
$P=|P|\otimes\mathbb{1}\otimes\sigma_{x}\otimes\mathbb{1}+\mathbb{1}\otimes|P|\otimes\mathbb{1}\otimes\sigma_{x},$ (18)
where we account for the possibility that the particles may be moving in different directions. This operator can also be expressed in block form, similar to the Hamiltonian in (16), where:
${\cal P}_{ab}=\sqrt{m}\begin{pmatrix}0&\sqrt{\epsilon+a\omega}&\sqrt{\epsilon+b\omega}&0\\ \sqrt{\epsilon+a\omega}&0&0&\sqrt{\epsilon+b\omega}\\ \sqrt{\epsilon+b\omega}&0&0&\sqrt{\epsilon+a\omega}\\ 0&\sqrt{\epsilon+b\omega}&\sqrt{\epsilon+a\omega}&0\end{pmatrix}.$ (19)
For a density matrix $\rho$ also expressed in diagonal blocks $\rho=\text{diag}\left(\rho_{11},\rho_{-11},\rho_{1-1},\rho_{-1-1}\right)$, where each $\rho_{ab}$ is a $4\times 4$ $X$-shaped matrix, the problem reduces to solving four blocks of differential equations for the six independent functions that constitute the elements of $\rho_{ab}$. Specifically, the Lindblad equation becomes:
$\partial_{t}\rho_{ab}+i\left[{\cal H}_{ab},\rho_{ab}\right]+\frac{\ell}{2}\left({\cal P}_{ab}^{2}\rho_{ab}+\rho_{ab}{\cal P}_{ab}^{2}-2{\cal P}_{ab}\rho_{ab}{\cal P}_{ab}\right)=0.$ (20)
Since the particles can move in any direction, there is a $25\%$ probability for the system to be in one of the four possible configurations of parameters $(a,b)=\{-1,1\}$. Therefore, the density matrix used to compute correlations must be averaged as:
$\bar{\rho}=0.25\sum_{a,b}\rho_{ab}.$ (21)
with elements
$\bar{\rho}_{ij}\doteq[\bar{\rho}]_{ij}=0.25\sum_{a,b}[\rho_{ab}]_{ij}$ (22)
To analyze this state, we consider solutions to the Lindblad equation Carrion:2024lap , assuming the initial condition:
$|\Psi\rangle=\sin\theta|01\rangle+\cos\theta|10\rangle,$ (23)
which represents an unentangled state when $\theta=0$ and a maximally entangled state (Bell state) when $\theta=\pi/4$. Thus, the parameter $\theta$ allows us to control the initial level of entanglement. The corresponding averaged matrix for this initial conditions is
$\bar{\rho}(t=0)=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&\cos^{2}(\theta)&\sin(\theta)\cos(\theta)&0\\ 0&\sin(\theta)\cos(\theta)&\sin^{2}(\theta)&0\\ 0&0&0&0\\ \end{array}\right).$ (28)
IV Quantum Correlations
Quantum spacetime is typically responsible for introducing dissipation in the form of decoherence in quantum systems (see, for instance, Petruzziello:2020wkd ). However, a unique signature of the deformation of the relativity principle is the presence of modified composition laws for energy and momentum. For single-particle states, decoherence emerges as a consequence of the modified evolution equation (10), as discussed in Arzano:2022nlo . However, for multiple particles, the introduction of interaction-like terms that depend on the momenta of each particle may introduce correlations that initially compete with decoherence effects. In this section, we will consider founr quantifiers of correlations: concurrence, as a measure of entanglement; the $l_{1}$-norm of coherence, quantum discord, and local quantum Fisher information.
IV.1 Concurrence
Entanglement is the most profound quantum correlation between systems and was first highlighted in the work by Einstein, Podolsky, and Rosen PhysRev.47.777 , which aimed to point out an incompleteness in the quantum description of the world. The modern definition of entanglement dates back to the works by Werner PhysRevA.40.4277 , where this concept was extended to mixed states. Today, entanglement plays a fundamental role in quantum cryptography PhysRevLett.67.661 , quantum computing PhysRevLett.69.2881 , and quantum teleportation PhysRevLett.70.1895 .
A useful quantifier of entanglement for a two-qubit state, described by an $X$-shaped matrix, is the concurrence. For a $4\times 4$ matrix, it reads Wootters:1997id ; Hill:1997pfa :
$\mathcal{C}=2\max\left\{|\bar{\rho}_{23}|-\sqrt{\bar{\rho}_{11}\bar{\rho}_{44}},|\bar{\rho}_{14}|-\sqrt{\bar{\rho}_{22}\bar{\rho}_{33}}|,0\right\},$ (29)
where the density matrix used is the averaged one described in (21). The system of differential equations (20) for the initial condition (23) cannot be solved analytically for arbitrary $\theta$, but we can describe the behavior of the concurrence for different parameters.
In Fig. 1, we observe the concurrence for different initial conditions. The qualitative behavior of the quantities discussed in this paper does not change significantly for different values of the parameter $\ell$. Only the scale in which they occur is modified. Therefore, without loss of generality, we consider $\ell$ with the same order of magnitude of the other parameters of the model. We see that a system of two particles can be prepared in an unentangled state ($\theta=0$) and become entangled over time due to the modified energy conservation in deformed relativity (black, solid curve). By increasing the angle to $\theta=0.05$, we see that although the state is initially partially entangled, decoherence effects reduce the concurrence to zero. However, the modified composition law effects are still strong enough to reentanglethe system before it becomes completely unentangled (blue, dash-dotted curve). For $\theta=\pi/12$, the system starts more entangled before experiencing a drop in this correlation (purple, dashed curve). This behavior repeats for the initial Bell state ($\theta=\pi/4$) (red, dotted curve). We considered unrealistically large values of the parameters for the sake of clarity of the figures, but this does not affect the main results. In fact, we are considering a value of the $\ell$-parameter of the same order of the other parameters in this plot, but in reality, it should be much smaller, since it is related to the Planck length scale of quantum gravity, which justifies the use of the Lindblad equation in this Markovian approximation.
Figure 1: Concurrence for different initial conditions determined by $\theta$. The parameters used are $\ell=1$, $m=1$, $\epsilon=1$, and $\omega=0.5$.
If we consider the product state determined by $\theta=0$, then, from Yang:2017llj , the time scale that governs the entanglement of this system is given by $T_{\text{ent}}=(\ell\Delta P^{2})^{-1}$, where $\Delta P$ is the standard quantum uncertainty of the momentum. If we average the uncertainty of the observable as $\Delta P=0.25\sum_{ab}\Delta P_{ab}$ and express the result by restoring the speed of light and Planck’s constant, we find:
$T_{\text{ent}}=\frac{\hbar E_{\text{QG}}}{2mc^{2}\epsilon},$ (30)
where $E_{\text{QG}}=\ell^{-1}$ is the quantum gravity energy scale that measures deviations from standard symmetries. Notice that this expression has an energy dependence similar to the decoherence time found in Arzano:2022nlo , $\tau_{D}\sim 2E_{\text{QG}}/p^{2}\sim E_{\text{QG}}/(mE)$. For the parameters used in Fig. 1, $T_{\text{ent}}$ is of the order of $0.5$ in natural units, which is consistent with Fig. 1. To understand this, let us suppose that we prepare such an unentangled superposition of two free particles. The uncertainty in the momentum of each particle translates into the time that it takes to observe the entanglement generation between them.
If the energy scale of the particles is $E$, then $T_{\text{ent}}E\sim\hbar E_{\text{QG}}/(mc^{2})$. For instance, consider a pair of ions Krutyanskiy:2022yoa with a mass of the order $m=3.7\times 10^{10}\,\text{eV}$, where the energy scale of electronic transitions is of the order $E\sim 1\,\text{eV}$, the Planck energy $E_{\text{QG}}=E_{\text{Planck}}=1.2\times 10^{28}\,\text{eV}$ and $\hbar=6.6\times 10^{-16}\text{eV}\cdot\text{s}$. We can estimate the entanglement time scale $T_{\text{ent}}=214\,\text{s}\approx 3.6\,\text{min}$. If an entangled pair of ions has approximately $100$ elementary particles and supposing that the quantum gravity energy scale grows linearly with the number of particles, the entanglement time becomes of the order of hours. This means that it would be necessary to preserve the quantum properties of this superposition of two particle system along this time scale in order to probe entanglement generation due to the uncertainty in their momenta. Besides that, if the quantum gravity scale $E_{\text{QG}}$ grows with the number of elementary particles, then the entanglement time scale grows too.
An interesting opportunity could be achieved when we consider long lived qubits, like trapped ions. For instance, in wang2020single it is reported coherence time exceeding one hour for a single qubit. Coherence time of system of qubits can serve as laboratory tests for deformations of the conservation of energy and deformed Hamiltonians.
IV.2 The $l_{1}$-Norm of Coherence
Coherence is an important ingredient in various areas of physics, such as quantum optics, quantum information, solid-state physics, and biology li2012witnessing . The first axiomatic approach to quantifying coherence was proposed in Aberg:2006xcv , and since then, several approaches have been developed. For a review, we refer the reader to Streltsov:2016iow . In this paper, we use the $l_{1}$-norm of coherence, introduced by Baumgratz:2013ecx , which is a quantifier based on matrix norms and can be straightforwardly applied to the type of density matrix we are analyzing.
For an $X$-shaped density matrix, it reads:
$C_{l_{1}}=\sum_{i\neq j}|\bar{\rho}_{ij}|=2(|\bar{\rho}_{14}|+|\bar{\rho}_{23}|),$ (31)
which simply measures the magnitude of the off-diagonal elements of $\bar{\rho}$.
In Fig. 2, we see that for $\theta<\pi/4$, the Hamiltonian creates correlations and initially dominates over decoherence effects, sharing the behavior of the concurrence discussed in the previous subsection. Over time, the Lindbladian terms eventually balance the deformed composition law contribution. Asymptotically, we observe a perfect balance between the creation of correlations due to the Hamiltonian and decoherence effects due to the Lindblad equation, which preserves the coherence of the system. Notably, for an initial Bell state ($\theta=\pi/4$), since the system begins maximally coherent, these two contributions balance each other from the outset. In fact, the open quantum spacetime also creates correlations in this system, as we will discuss in Section V, meaning that while it dissipates entanglement, it helps maintain other correlations between the particles.
Figure 2: The $l_{1}$-norm of coherence for different initial conditions determined by $\theta$. The parameters used are $\ell=1$, $m=1$, $\epsilon=1$, and $\omega=0.5$.
This behavior is not entirely new in condensed matter systems. For instance, a similar phenomenon can be observed in Carrion:2024lap , which considers a spin-$1/2$ Ising-$XYZ$ chain model. There, one also observes the disappearance and resurgence of entanglement and coherence, with the stabilization of quantum correlations over time.
IV.3 Quantum Discord
Quantum discord serves as a quantifier of quantum correlations beyond entanglement. By calculating it, we can determine whether the balance between the Hamiltonian and the Lindbladian dissipates all quantum correlations or only entanglement. The definition of quantum discord uses the concept of mutual information between two random variables Streltsov2015 ; PhysRevLett.88.017901 . For a system consisting of two subsystems, say $A$ and $B$ (e.g., the two particles considered in this paper), the density matrix is written as $\rho^{AB}$, which in our case is $\bar{\rho}$. The total mutual information is given by:
$I(\rho^{AB})=S(\rho^{A})+S(\rho^{B})-S(\rho^{AB}),$ (32)
where $S(\rho)=-\text{Tr}[\rho\log_{2}\rho]$ is the von Neumann entropy, which measures the uncertainty of a quantum state $\rho$.
The quantity $\rho^{A}$ is the reduced density operator of subsystem $A$, defined as:
$\rho^{A}=\text{Tr}_{B}[\rho^{AB}],$ (33)
where $\text{Tr}_{B}$ is the partial trace over subsystem $B$, given by:
$\text{Tr}_{B}[\rho^{AB}]=\sum_{k=1}^{d}({\mathbb{1}}_{A}\otimes\langle k_{B}|)\rho^{AB}({\mathbb{1}}_{A}\otimes|k\rangle_{B}),$ (34)
where $d$ is the dimension of subsystem $B$, ${\mathbb{1}}_{A}$ is the identity operator of subsystem $A$ (in our case, $\text{diag}(1,1)$), and $\{|k\rangle_{B}\}$ is an orthonormal basis of subsystem $B$ maziero2017computing ; Watrous:2018zil . In our case, $d=2$, and we consider the computational basis $\{|k\rangle_{B}\}=\{|0\rangle,|1\rangle\}$. For a $4\times 4$ matrix with elements $[\rho^{AB}]_{ij}=\rho_{ij}$, we have:
$\rho^{A}=\text{Tr}_{B}(\rho^{AB})=\begin{pmatrix}\rho_{11}+\rho_{33}&\rho_{12}+\rho_{34}\\ \rho_{21}+\rho_{43}&\rho_{22}+\rho_{44}\end{pmatrix},$ (35)
$\rho^{B}=\text{Tr}_{A}(\rho^{AB})=\begin{pmatrix}\rho_{11}+\rho_{22}&\rho_{13}+\rho_{24}\\ \rho_{31}+\rho_{42}&\rho_{33}+\rho_{44}\end{pmatrix}.$ (36)
This procedure provides a way to derive the state of $A$ by tracing out the degrees of freedom of $B$, and vice versa. It ensures the correct measurement statistics for measurements made on subsystems $A$ and $B$ Nielsen:2012yss .
For a classical system, there is an equivalent way to define mutual information. The Shannon entropy is the classical counterpart of the von Neumann entropy. Consider a classical system described by two random variables $X$ and $Y$, where $p_{x}$ is the probability that $X$ takes the value $x$, $p_{Y}$ is the probability that $Y$ takes the value $y$, and $p_{x,y}$ is the probability that a measurement of $(X,Y)$ yields $(x,y)$. The Shannon entropy for $X$ is given by $H(X)=-\sum_{x}p_{x}\log_{2}p_{x}$.
From Bayes’ rule, the probability that $X$ takes the value $x$, given that $Y$ takes the value $y$, is $p_{x|y}=p_{xy}/p_{y}$. This allows us to write:
$H(X,Y)=-\left(\sum_{y}p_{y}\log_{2}p_{y}\right)\left(\sum_{x}p_{x|y}\right)$ (37)
$-\sum_{y}p_{y}\left(\sum_{x}p_{x|y}\log_{2}p_{x|y}\right)=H(Y)+H(X|Y),$
where $H(X|Y)=\sum_{y}H(X|y)$ is the conditional entropy, representing the amount of information needed to describe the outcome of $X$ given the outcome of $Y$. Thus, there are two equivalent ways, $I$ and $J$, to describe mutual information for a classical system:
$I(X,Y)$ $=H(X)+H(Y)-H(X,Y),$ (38)
$J(X,Y)$ $=H(X)-H(X|Y).$ (39)
A quantum analog of the conditional entropy for a bipartite system $\rho^{AB}$ was introduced by Ollivier and Zurek PhysRevLett.88.017901 :
$S(A|\{\Pi_{i}^{B}\})=\sum_{i}p_{i}S(\rho_{i}^{A}),$ (40)
where $\{\Pi_{i}^{B}\}$ are orthogonal measurement operators on subsystem $B$, and $p_{i}=\text{Tr}[\Pi_{i}^{B}\rho^{AB}]$ is the probability of obtaining outcome $i$. Additionally, $\rho_{i}^{A}=\text{Tr}_{B}[\Pi_{i}^{B}\rho^{AB}]/p_{i}$ is the post-measurement state of subsystem $A$. Thus, the quantum version of $J$ defined in (39) is:
$J(\rho^{AB})_{\{\Pi_{i}^{B}\}}=S(\rho^{A})-S(A|\{\Pi_{i}^{B}\}),$ (41)
which depends on the measurement operator $\Pi_{i}^{B}$ and represents the amount of information gained about $A$ by measuring $B$.
The quantum discord is the difference between the expressions (38) and (41), minimized over all von Neumann measurements  [^1] A general quantum measurement is described by a collection $\{E_{i}\}$ of measurement operators that satisfy the completeness equation: $\sum_{i}E^{\dagger}_{i}E_{i}={\mathbb{1}}$. For a projective measurement, the operators $E_{i}$ are orthogonal projectors: $E_{i}E_{j}=\delta_{ij}E_{i}$. A von Neumann measurement is a special type of projective measurement where the measurement operators $E_{i}$ are orthogonal projectors with rank one (i.e., their image is a column matrix).. In our case, the optimal measurement operator is $\{\Pi_{i}^{B}\}=\{|0\rangle\langle 0|,|1\rangle\langle 1|\}$.
Therefore, the quantum discord is given by:
$D=S(\rho^{B})-S(\rho^{AB})+p_{0}S(\rho_{0}^{A})+p_{1}S(\rho_{1}^{A}),$ (42)
where $p_{0}=\text{Tr}[|0\rangle\langle 0|\rho^{AB}]$, $p_{1}=\text{Tr}[|1\rangle\langle 1|\rho^{AB}]$, $\rho_{0}^{A}=\text{Tr}_{B}[|0\rangle\langle 0|\rho^{AB}]/p_{0}$, $\rho_{1}^{A}=\text{Tr}_{B}[|1\rangle\langle 1|\rho^{AB}]/p_{1}$, and $\rho^{AB}=\bar{\rho}$, as given by (21). In Appendix A, we express this quantity for any $4\times 4$ density matrix.
We depict this quantifier for different initial conditions in Fig. 3. As can be seen, the quantum discord behaves similarly to the $l_{1}$-norm of coherence. It also stabilizes over very long times, indicating that while entanglement degrades over time, other quantum correlations between the two particles persist.
Figure 3: Quantum discord for different initial conditions determined by $\theta$.
We also depict the behavior of the three correlations discussed in this paper for an initially unentangled state in Fig. 4. It is evident that there is a hierarchy between coherence, quantum discord and entanglement.
Figure 4: Comparison of the different quantifiers for coherence, discord and entanglement studied in the paper for $\theta=0$. The red/dotted curve represents the $l_{1}$-norm of coherence. The blue/dash-dotted curve represents quantum discord. The black/solid curve represents concurrence.
IV.4 Local Quantum Fisher Information
The Quantum Fisher Information (QFI) is a fundamental concept in quantum estimation theory, which measures how much information a quantum state carries about a parameter to be estimated. It is related to the maximum precision with which we can estimate a parameter through the quantum Cramér-Rao bound slaoui2019comparative ; Silva:2025vmu ; kim2018characterizing ; Dahbi:2022uyi ; Lian:2021xjw ; Yurischev:2023jrd .
In many applications of quantum mechanics and quantum information, we aim to estimate a parameter $\phi$ that affects a quantum state $\rho(\phi)$. This parameter can represent, for example: a magnetic field applied to a quantum system, an evolution time of a quantum state or a phase shift in an optical system. The idea is to measure a quantum system to obtain as much information as possible about $\phi$.
The Quantum Fisher Information is a generalization of classical Fisher Information to quantum states. The formal definition is:
$F_{Q}(\phi)=Tr\left(\rho(\phi)L_{\phi}^{2}\right)\,,$ (43)
where $L_{\phi}$ is the quantum score operator, also called the Helstrom operator, defined implicitly by the Lyapunov equation:
$\frac{d\rho(\phi)}{d\phi}=\frac{1}{2}(\rho(\phi)L_{\phi}+L_{\phi}\rho(\phi)).$ (44)
The operator $L_{\phi}$ represents the best choice of observable for estimating the parameter $\phi$.
The QFI is widely used in various areas of physics and quantum information, such as: quantum metrology, for improving the precision of sensitive measurements, such as interferometry and quantum sensors; quantum communication, for optimizing the transmission and detection of signals encoded in quantum states; quantum computing, for optimizing algorithms based on phase estimation and entangled states, quantum thermodynamics, for understanding the efficiency of quantum heat engines.
In particular, when the parametric states $\rho_{\phi}$ can be obtained from an initial probe state $\rho$ subjected to a unitary transformation $U_{\phi}=e^{iH\phi}$ dependent on $\phi$ and generated by a Hermitian operator $H$, that is, $\rho_{\phi}=U_{\phi}^{\dagger}\rho U_{\phi}$. In this case, the Quantum Fisher Information $F(\rho_{\phi})$, which we denote by $F(\rho,H)$, is given by:
$F(\rho,H)=\sum_{i\neq j}\frac{(p_{i}-p_{j})^{2}}{p_{i}+p_{j}}|\langle\psi_{i}|H|\psi_{j}\rangle|^{2}\,.$ (45)
Here, we use the spectral decomposition of $\rho$, that is,
$\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|,$ (46)
with $p_{i}\geq 0$ and $\sum_{i}p_{i}=1$  . The main terms here are:
  • $p_{i}$: eigenvalues of the density matrix $\rho$, obtained from its spectral decomposition $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$.
  • $|\psi_{i}\rangle$: eigenvectors corresponding to $\rho$.
  • $H$: Hermitian operator associated with the unitary evolution of the system.
  • $\frac{(p_{i}-p_{j})^{2}}{p_{i}+p_{j}}$: weight that quantifies the difference between the eigenvalues of the quantum state $\rho$.
  • $|\langle\psi_{i}|H|\psi_{j}\rangle|^{2}$: matrix element of $H$ between the eigenvectors $|\psi_{i}\rangle$ and $|\psi_{j}\rangle$.
The important point here is that the QFI is independent of $\phi$.
Now, consider a bipartite quantum state $\rho_{AB}$ of dimension $2\times d$ in the Hilbert space $H=H_{A}\otimes H_{B}$. We assume that the dynamics of the first part is governed by the local phase shift transformation $e^{-i\phi H_{A}}$, where $H_{A}\equiv H_{a}\otimes I_{B}$ is the local Hamiltonian. In this case, the QFI reduces to the Local Quantum Fisher Information (LQFI), given by:
$F(\rho,H_{A})=Tr(\rho H_{A}^{2})-\sum_{i\neq j}\frac{2p_{i}p_{j}}{p_{i}+p_{j}}|\langle\psi_{i}|H_{A}|\psi_{j}\rangle|^{2}\,.$ (47)
The Local Quantum Fisher Information (LQFI) was introduced to deal with quantum correlations of the discord type. This quantifier allows a deeper understanding of how quantum correlations influence metrological precision. It has desirable properties that any good quantifier of quantum correlations should satisfy:
  • It is non-negative and vanishes for bipartite states without discord (classically correlated states).
  • It is invariant under local unitary operations.
  • For pure states, it coincides with the geometric discord.
The goal of choosing the local Hamiltonian was to capture the quantum correlations between the subsystems. If the system is purely classical, then the LQFI will be zero, as a local observer does not detect quantum effects.
The quantification of quantum correlations in terms of the Local Quantum Fisher Information $Q(\rho)$ is defined as the minimization of the QFI over all local Hamiltonians $H_{A}$ acting on part $A$:
$Q(\rho)=\min_{H_{A}}F(\rho,H_{A})\,.$ (48)
The general form of a local Hamiltonian is
$H_{A}=\vec{\sigma}\cdot\vec{r},$
with $|\vec{r}|=1$ and $\vec{\sigma}=(\sigma_{x}\equiv\sigma_{1},\sigma_{y}\equiv\sigma_{2},\sigma_{z}\equiv\sigma_{3})$ being the usual Pauli matrices.
It can be verified that
$Tr(\rho H_{A}^{2})=1,$
and the second term in equation (2) can be rewritten as:
$\sum_{i\neq j}\frac{2p_{i}p_{j}}{p_{i}+p_{j}}|\langle\psi_{i}|H_{A}|\psi_{j}\rangle|^{2}=\vec{r}^{\dagger}M\vec{r},$
where the elements of the symmetric matrix $M$ of dimension $3\times 3$ are given by:
$M_{lk}=\sum_{i\neq j}\frac{2p_{i}p_{j}}{p_{i}+p_{j}}\langle\psi_{i}|\sigma_{l}\otimes I_{B}|\psi_{j}\rangle\langle\psi_{j}|\sigma_{k}\otimes I_{B}|\psi_{i}\rangle.$ (49)
To minimize $F(\rho,H_{A})$, it is necessary to maximize the quantity $\vec{r}^{\dagger}M\vec{r}$ over all unit vectors $\vec{r}$. The maximum value of this expression corresponds to the largest eigenvalue of the matrix $M$.
Therefore, the minimum value of the Local Quantum Fisher Information $Q(\rho)$ is given by:
$Q(\rho)=1-\lambda_{\max}(M),\quad(8)$
where $\lambda_{\max}$ represents the largest eigenvalue of the symmetric matrix $M$ defined in equation (49).
Figure 5: The local quantum Fisher information $(LQFI)$ as a function of time $t$ for different initial quantum states. The black curve represents the $LQFI$ for the non-entangled initial state $(\theta=0)$, while the red curve corresponds to the entangled initial state $(\theta=\pi/4)$. The parameters used are $\ell=1$, $m=1$, $\epsilon=1$, and $w=0.5$.
Fig. 5 illustrate the behavior of $LQFI$ as a function of time $t$, comparing the $LQFI$ for the initial state $(\theta=0)$ with that of the initial state $(\theta=\pi/4)$. This figure reveals that the quantum correlations embedded in the $LQFI$ for the initially non-entangled state are more robust than those for the initially entangled state. This occurs because entangled states are more sensitive to decoherence than separable states, causing the $LQFI$ to degrade more rapidly.
V The Role of the Different Planck Scale Contributions
To better understand the effect of the deformed Hamiltonian in creating correlations and the Lindbladians in destroying them, we introduce two deformation parameters, $\ell_{H}$ and $\ell_{L}$, to control these contributions. The Hamiltonian is given by:
${\cal H}$ $=\epsilon(\mathbb{1})^{4}+\dfrac{\omega}{2}\left(\sigma_{z}\otimes\mathbb{1}\otimes\mathbb{1}\otimes\mathbb{1}+\mathbb{1}\otimes\sigma_{z}\otimes\mathbb{1}\otimes\mathbb{1}\right)$
$\quad+\ell_{H}|P|\otimes|P|\otimes\sigma_{x}\otimes\sigma_{x},$ (50)
and the Lindblad equation is given by:
$\partial_{t}\rho_{ab}+i\left[{\cal H}_{ab},\rho_{ab}\right]+\frac{\ell_{L}}{2}\left({\cal P}_{ab}^{2}\rho_{ab}+\rho_{ab}{\cal P}_{ab}^{2}-2{\cal P}_{ab}\rho_{ab}{\cal P}_{ab}\right)=0.$ (51)
V.1 Lindbladian Weaker than Hamiltonian ($\ell_{L}\ll\ell_{H}$)
The effect of the different contributions to the Lindblad equation can be seen in Fig. 6, where we assume different values for the deformation parameters. The dissipation effect is described by $\ell_{L}=0.005$, while the deformed Hamiltonian is described by $\ell_{H}=1$. For an initially unentangled state ($\theta=0$, black curve), we observe a dynamic phenomenon of alternating sudden death and sudden birth of quantum entanglement between the particles, whose amplitude is progressively attenuated by the dissipation introduced by the Lindblad term. In contrast, for a partially entangled initial state ($\theta=\pi/6$, blue curve), the concurrence drops monotonically until it vanishes.
Figure 6: Concurrence for different initial conditions determined by $\theta$. The parameters used are $\ell_{H}=1$, $\ell_{L}=0.005$, $m=1$, $\epsilon=1$, and $\omega=0.5$.
Similarly, the $l_{1}$-norm of coherence and quantum discord exhibit oscillatory behavior, as shown in Figs. 7 and 8. For an initially unentangled state ($\theta=0$, black curve) and for the partially entangled state ($\theta=\pi/6$, red curve), the coherence oscillates but tends to stabilize over time, as observed in the case where $\ell_{H}=\ell_{L}$. To further study the role of these contributions, let us consider the opposite case. If $\ell_{L}$ were zero, the system’s concurrence and $l_{1}$-norm of coherence would oscillate without dissipation.
Figure 7: The $l_{1}$-norm of coherence for different initial conditions determined by $\theta$. The parameters used are $\ell_{H}=1$, $\ell_{L}=0.005$, $m=1$, $\epsilon=1$, and $\omega=0.5$.
Figure 8: Quantum discord for different initial conditions determined by $\theta$. The parameters used are $\ell_{H}=1$, $\ell_{L}=0.005$, $m=1$, $\epsilon=1$, and $\omega=0.5$.
V.2 Hamiltonian Weaker than Lindbladian ($\ell_{H}\ll\ell_{L}$)
In this case, when the system starts in an unentangled state, the deformed composition law contribution is not strong enough to produce non-zero concurrence. For an initially entangled state, the concurrence evolves similarly to the behavior shown in Fig. 1 for the same set of parameters.
Regarding the $l_{1}$-norm of coherence and quantum discord, the system behaves similarly to Figs. 2 and 3. This provides evidence that the open quantum system is actively creating and maintaining superpositions between the particles. Although it dissipates entanglement, it helps create and preserve classical and quantum correlations beyond entanglement. The reason for this behavior is in the fact that the Lindbladian ${\cal P}_{ab}$ is not diagonal due to the superposition of momenta in this model, as can be seen in (19). This property is responsible for giving a non-null contribution from the open system represented by the quantum spacetime and creating superpositions in this model.
VI Discussion
We analyzed how modifications to the Hamiltonian of a system of particles, arising from deformations of the relativistic symmetries, influence quantum correlations. Specifically, we constructed a model describing two particles in a superposition of energies and momenta, corresponding to a system of four qubits. This system exhibits intriguing properties, such as the emergence of entanglement due to the modified Hamiltonian and the role of the quantum spacetime as an environment that dissipates entanglement while simultaneously creating and preserving other classical and quantum correlations between the particles. This behavior was observed when we individually examined the contributions of the quantum gravity parameters from the deformed energy composition law (Hamiltonian) and from the Lindblad equation.
The unique ability of the deformed Hamiltonian to generate entanglement between the particles in the system is a distinctive effect of a deformation of relativistic symmetries. If coherence is maintained during the time scale of the entanglement time described in this paper, then systems of qubits could be considered to test deformations of the composition of energy due to quantum gravity.
Appendix A Quantum Discord for Any $4\times 4$ Density Matrix
Following the steps described in Section IV.3, the quantum discord for a bipartite system (represented by a $4\times 4$ density matrix), using the computational basis for optimal von Neumann measurements, is given by:
$D$ $=\frac{1}{2\log 2}\Bigg{(}-\frac{2(\rho_{22}+\rho_{44})\left(\rho_{11}\log\frac{\rho_{11}}{\rho_{11}+\rho_{22}}+\rho_{22}\log\frac{\rho_{22}}{\rho_{11}+\rho_{22}}\right)}{\rho_{11}+\rho_{22}}$
$\quad-2(\rho_{11}+\rho_{22})\log(\rho_{11}+\rho_{22})$
$\quad+\left(\rho_{22}+\rho_{33}-\sqrt{(\rho_{22}-\rho_{33})^{2}+4|\rho_{23}|^{2}}\right)\log\frac{1}{2}\left(\rho_{22}+\rho_{33}-\sqrt{(\rho_{22}-\rho_{33})^{2}+4|\rho_{23}|^{2}}\right)$
$\quad+\left(\rho_{22}+\rho_{33}+\sqrt{(\rho_{22}-\rho_{33})^{2}+4|\rho_{23}|^{2}}\right)\log\frac{1}{2}\left(\rho_{22}+\rho_{33}+\sqrt{(\rho_{22}-\rho_{33})^{2}+4|\rho_{23}|^{2}}\right)$
$\quad-\frac{2(\rho_{11}+\rho_{33})\left(\rho_{33}\log\frac{\rho_{33}}{\rho_{33}+\rho_{44}}+\rho_{44}\log\frac{\rho_{44}}{\rho_{33}+\rho_{44}}\right)}{\rho_{33}+\rho_{44}}$
$\quad-2(\rho_{33}+\rho_{44})\log(\rho_{33}+\rho_{44})$
$\quad+\left(\rho_{11}+\rho_{44}-\sqrt{(\rho_{11}-\rho_{44})^{2}+4|\rho_{14}|^{2}}\right)\log\frac{1}{2}\left(\rho_{11}+\rho_{44}-\sqrt{(\rho_{11}-\rho_{44})^{2}+4|\rho_{14}|^{2}}\right)$
$\quad+\left(\rho_{11}+\rho_{44}+\sqrt{(\rho_{11}-\rho_{44})^{2}+4|\rho_{14}|^{2}}\right)\log\frac{1}{2}\left(\rho_{11}+\rho_{44}+\sqrt{(\rho_{11}-\rho_{44})^{2}+4|\rho_{14}|^{2}}\right)\Bigg{)}.$ (52)
Acknowledgments
I. P. L. would like to thank Adriano Braga Barreto for the insightful discussions. I. P. L. was partially supported by the National Council for Scientific and Technological Development - CNPq, grant 312547/2023-4. G. V. thanks Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES) - Finance Code 001 and the National Council for Scientific and Technological Development - CNPq, grant 140335/2022-6 for financial support. V. B. B. was partially supported by the National Council for Scientific and Technological Development- CNPq, grant No. 307211/2020 7.
The authors would like to acknowledge networking support by the COST Action BridgeQG (CA23130), supported by COST (European Cooperation in Science and Technology).